# A Survey on Computational Solutions for Reconstructing Complete Objects by Reassembling Their Fractured Parts

### v1

@@ 7.1 Bone Reduction
Deep learning in bone reduction. In the literature, the deep learning models were trained to segment the cortical surface of bone fragments [234] and fractured bones [254], identify and localize bone fractures [240], and compute bone features [214, 234] for template matching. Recently, a significant study [231] introduced an innovative bidirectional framework for automated pelvic fracture surgical planning, which integrates fracture simulation with structural restoration. The fracture simulation module specifically considered patient-specific pelvic anatomy, bone density variations, and fracture stochasticity to generate diverse fracture scenarios from intact pelvises. For the creation of training datasets that incorporate actual fractures, images of the fractured side and contralateral healthy side were utilized as input and ground-truth data, respectively. Accordingly, being trained on both simulated and real fractured bones, the proposed deep structural restoration network reconstructs the bone’s pre-fracture state, subsequently serving as the template for matching the input bone fragments.
7.2 Fractured Thin-shell Object Reassembly
In the context of 3D reassembly tasks, a thin-shell object refers to a structure with a thickness that is negligible compared to its other dimensions. This implies that the object’s surface geometry, rather than its volume, primarily defines its shape. Thin-shell objects present significant challenges during reassembly due to their intricate nature, making the matching and alignment of fragments more complex.
In forensic skull reassembly, the skull fragments are considered thin shell objects because their thickness is minimal compared to their overall size. This requires a precise matching of the fracture regions and intact surfaces to accurately reconstruct the original shape. Yin et al. [200] and Wei et al. [201] model the set of skull fragments as a collection of surfaces. They detect the boundary loops of the skull fragments as feature curves and then compute feature descriptors on these curves. The feature curves are used for pair-wise matching and alignment computation. However, Zhang et al. [189] note that due to the thickness of the skull fragments, the feature curves can form multiple loops and may be ambiguous in a single fragment. To match multiple feature curves, they formulate it as the Largest Common Point-set (LCP) problem. They developed a voting algorithm using 4-point congruent sets extraction to solve this LCP problem.
@@ 7.2 Fractured Thin-shell Object Reassembly
Another relevant scenario is archaeological artifact reassembly such as pottery [87, 255], where broken fragments are often thin. Cohen et al. [87] first digitized the excavated fragments using a Konica 3D scanner (See Figure 17). They treated the scanned fragments as surfaces and computed the parabolic contours on them. Finally, they used affine-invariant 3D moments to match the contours and align the fragments. Their system shows a magnitude improvement in running time compared to manual solutions.
Angelo et al. [256] proposed a recognition pipeline to determine the regions of the 3D scanned fragments into the rim, internal/external wall, and base. To reassemble a complete pottery, they designed various descriptors for different regions of the fragments. These descriptors help find correspondences and align the fragments. Fu et al. [257] take curvature and torsion as characteristic descriptors of the curve and search for the candidate matching by the derivative dynamic time warping (DDTW). They then design a voting mechanism and ICP method to align the fragments. Wang et al. [24] observed that due to the thin shell nature of pottery fragments, thin side strip regions often contain many repetitive fractured patterns and lack sufficient unique geometric features. They provide a novel contour-based solution to solve batch matching and a boundary-aware ICP method to tackle fine registration. They devised an intrinsic shape descriptor to encode a 2D contour to fuse the front and back scans. After each fragment’s front and back scans are paired and initially aligned, to address the issue of small overlap and the lack of salient textures and features in the small overlapping region of the two scans, they propose a boundary-based ICP method to ensure robust ICP iteration. Their approach demonstrated the ability to scan and reconstruct over 700 fragments in 10 working hours, achieving high reconstruction accuracy.
7.3 Wall Fragment Reassembly
@@
Figure 18: The top row shows projects of restoring wall fragments. The bottom row shows projects of restoring building fragments. (Top-left). Stanford’s Forma Urbis Romae project [258]. (Top-right) Princeton’s . Theran wall paintings project [22]. (Bottom-left). Restoration of the Octagon monument in Ephesos [50]. (Bottom-right) Reconstruction of the Arch of Titus at the Circus Maximus in Rome [206].
A special class of thin-shell fragments are wall fragments, in which the complete objects is flat. The first well-known effort is Stanford’s Forma Urbis Romae project [258] (Figure 18 (Top-Left)). They scanned 1186 fragments that are physically available and, according to estimates, account for about 20% of all fragments. As a significant portion of the fragments are missing, they only report dozens of potential matches [79]. The methods to compute use methods include boundary incision matching, which employ surface textual features, wall feature matching, which is based on location prediction using a underlying map, and multivariate clustering. Multivariate clustering finds fragment pairs that may not have explicitly corresponding incisions, but may share other common characteristics including fragment thickness, marble veining direction, axial direction of the incised architecture, presence and orientation of slab edges, and the back surface condition of the fragments (rough, smooth, or unpreserved.
@@
Another category of notable projects are efforts to reconstruct and match Theran wall paintings [259, 260, 22, 261, 140, 262, 144], including a series of work at Princeton University on reconstruction [22], matching [260, 261, 140], and analysis of fragment boundary patterns for fragment pattern simulation [262]. Moreover, [144] presents a generic algorithm for reassembling wall paintings. One major difference between wall fragments and fragments of bones and thin-shells is that the number of wall fragments is much larger. In other words, it is important to develop computational solutions to fractured object reassembly. Another consequence of having a large number of fragments is that the global consistency constraint [140, 144] becomes important for multi-fragment matching.
One relevant scenario is the restoration of ancient papyrus fragments, which share the common characteristics of being flat. [263] addresses the reassembly of these fragments using a convolutional neural network in conjunction with a random forest classifier.
7.4 Fractured Building Reassembly
Another special class of fragments are fragments of architectural buildings. In this case, virtual restoration is necessary as it is very difficult to move the fragments. Similar to the setting of wall fragments, the number of fragments is much larger than in the case of bone fragments or thin shell fragments. Moreover, the percentages of missing fragments in the building setting, due to the open environment, are even higher than those in the wall painting setting. In this case, solving the fractured object reassembly problem strongly rely prior knowledge of the underlying complete object. Moreover, user input, i.e., semi-automatic approaches, are necessary to obtain good results.
One example is the project to reconstruct the Octagon monument in Ephesos [50]. The project used a semi-automatic approach to virtually restore the original monument from 130 fragments (about 20% of the total number of fragments). This restoration is heavily based on prior knowledge of the monument. Another relevant project is the reconstruction of the Arch of Titus at the Circus Maximus in Rome [206]. Again, many fragments are missing, and the virtual reconstruction relies on knowledge about the underlying symmetric structure of the Arch. Other notable efforts include the restorations of mixtilinear buildings in Villa Adriana [264], Macedonian Tombs [265], the Maritime Theatre at Hadrian’s Villa [266], and a project on restoring damaged column drums  [267].
@@ 7.5 Fractured Object Reassembly in Science
There are many scientific applications for fractured object reassembly. One such example is docking of proteins and DNA Sequencing. In [268], in which mitochondrial genes from the unicellular eukaryote Diplonema are systematically fragmented into small pieces that are encoded on separate chromosomes and individually transcribed. The reassembly problem is to concatenate them into contiguous messenger RNA molecules. Another biological application is protein docking [269, 270], where the goal is to find a rigid and non-rigid transformation of one 3D protein to interact with another protein.
Another application is in paleontology, in which we can reconstruct fossils from fragments precisely in 3D, to allow 3D analysis of the whole shape. A semi-automatic approach was introduced in [271] to reconstruct an extinct bird. It is desirable to develop automatic approaches for more accurate and faster reconstructions. [272] studied virtual reconstruction of huminin crania to understand human evolution. [273] introduced an interactive approach in Blender to restore a dorsal rib and frontal bone of a small-bodied Jurassic-age armored dinosaur from Africa, The digital restoration engaged all modalities of deformation (translation, rotation, scaling, distortion) and reconstruction (fracture infilling, adding missing bone, surface smoothing).
@@ 7.6 Part Assembly
Part assembly refers to the process of assembling components that are decomposed in a semantically meaningful way, such as furniture components like table legs[190] and mechanical parts like screws[275]. Although closely related to the fracture assembly problem, part assembly introduces several key differences. It generally assumes that each part has an independent semantic function and that each component is complete and functional on its own, unlike the fractured pieces in fracture assembly. Moreover, part assembly allows for duplicated components, and solutions are often non-unique. Unlike fracture assembly, where local or global geometry provides significant clues, part assembly relies more on semantic or functional understanding of the individual parts and the assembled object. However, despite these differences, we can identify valuable connections between the two problems, and both fields can benefit from exploring the shared foundations of these challenges.
@@
Figure 21: Datasets that are suitable for deep learning research in fractured object reassembly. (Top-Left) The real Fantastic Breaks dataset [295] which combines both isolated fragments (top) and paired complete objects and their fragments (bottom). (Top-Right) A real dataset of fragments with ceramic material [24] is captured using the system shown here. (Bottom) A synthetic dataset created by performing fracture simulation on a diverse set of objects [145].
@@
8 Datasets and Software Packages
We first discuss existing datasets in Section 8.1. We then discuss existing software packages 8.2.
@@ 8.1 Datasets
Machine learning approaches have shown great advancements in the task of reassembly of fragments, but are highly dependent on large-scale datasets. However, the first data sets used in fragment reassembly research are primarily small-scale, often manually created [52, 298, 258] or derived from automated semantic segmentation [293, 190].
Figure 22: Illustrations of future directions (Left) Representations and algorithms for matching complete template shapes with large geometric and topological variations and end-to-end learning of the entire reassembly pipeline. (Middle) Leveraging foundation models to solve sub-tasks such as segmenting fractured surfaces, matching fragments, predicting complete template shapes from input fragments, and matching complete template shapes and fragments. (Right) Learning to generate synthetic datasets of 3D fragments and and doing so jointly with learning the reassembly network.
@@
There are two recent efforts which have focused on scanning fragmented objects at scale. The first example is Fantastic breaks [295], which include a small collection of objects with broken fragments and more objects that were broken into fragments by the authors (see Figure 21(Left)). Both fragments and complete objects (if they exist) are scanned. Another example is [24], which developed a system to batch scan fragments of objects with ceramic material (see Figure 21(Right)). This dataset has ground-truth complete objects for evaluation. In addition, the scale of these two datasets makes it possible to develop deep neural networks for reassembly. However, the number of categories in these two datasets is still small, presenting generalization issues in novel categories.
@@
In addition to real data sets, Breaking Bad [145] is a synthetic data developed by running a simple fracture simulation procedure to generate fragments of shape collections that undergo large geometric and topological variations ((see Figure 21(Right)). This dataset is interesting for fractured object reassembly because the underlying objects belong to diverse object categories. For example, template-based reassembling techniques can be developed by learning a shape prior of the complete objects in this dataset. Just as the role of ShapeNet [299] in boosting 3D deep learning rich, we envision that research progress on algorithmic fractured object reassembly will be accelerated with Breaking Bad and similar datasets. However, as the fragments are generated using non-physics based fracture simulation. The question is how the reassembly pipelines derived from such synthetic datasets generalize to real data.
8.2 Software Packages

### v2

@@ 7.1 Bone Reduction
Deep learning models were trained to segment the cortical surface of bone fragments [240] and fractured bones [258], identify and localize bone fractures [244], and compute bone features [218, 240] for template matching. A notable advancement [237] introduced a bidirectional framework combining fracture simulation and structural restoration, where the simulation considers patient-specific anatomy and bone properties to generate training data. This approach uses both simulated scenarios and real fracture data, where contralateral healthy sides provide ground truth, to train a network that reconstructs pre-fracture states for template matching.
7.2 Fractured Cultural Heritage Reassembly
Cultural heritage restoration represents a primary application domain for computational reassembly techniques. Reconstructing damaged artifacts preserves historical objects and provides insights into past civilizations and artistic practices. We examine four categories of heritage objects: thin-shell artifacts like pottery, stone fragments from sculptures and buildings, planar objects such as maps and inscribed tablets, and larger-scale archaeological site reconstructions. Each category presents distinct geometric challenges - from pottery’s axial symmetry to stone’s complex fracture patterns - requiring specialized computational approaches that build upon core reassembly techniques discussed earlier.
@@ 7.2 Fractured Cultural Heritage Reassembly
Thin-shell objects. The reassembly of thin-shell archaeological artifacts, particularly pottery [90, 259] and ceramics [32, 227], presents unique challenges due to their distinctive geometric properties. These objects are characterized by surfaces where thickness is negligible compared to other dimensions, requiring specialized computational approaches. Early work by Cohen et al. [90] established a foundational pipeline using digitized fragments from 3D scans, computing parabolic contours on surfaces, and employing affine-invariant 3D moments for contour matching and alignment. This approach demonstrated significant speed improvements over manual methods. For ceramics specifically, researchers have leveraged their inherent axial symmetry properties. Karasik and Smilansky [32] developed efficient scanning protocols specifically for ceramic fragments, while Hong et al. [227] introduced an incremental 3D reassembly method that exploits axial symmetry for more robust matching.
Subsequent research enhanced the reconstruction process through various specialized techniques. [260] developed a recognition system that segments fragments into distinct regions (rim, walls, and base), designing region-specific descriptors to guide the matching process. [261] introduced a curve-based matching approach using curvature and torsion descriptors combined with derivative dynamic time warping (DDTW), incorporating voting mechanisms and ICP refinement. A recent breakthrough [24] addressed a critical challenge in thin-shell reassembly - the lack of distinctive geometric features inside regions. Their method combines contour-based batch matching with boundary-aware ICP registration, achieving remarkable efficiency by processing over 700 fragments within 10 hours while maintaining high accuracy. This progression of research demonstrates the evolution from basic geometric matching to sophisticated systems that can handle large-scale reconstruction tasks.
@@
Figure 19: The top row shows projects of restoring planar fragments. The bottom row shows projects of restoring building fragments. (Top-left). Stanford’s Forma Urbis Romae project [262]. (Top-right) Princeton’s . Theran wall paintings project [22]. (Bottom-left). Restoration of the Octagon monument in Ephesos [51]. (Bottom-right) Reconstruction of the Arch of Titus at the Circus Maximus in Rome [210].
Planar heritage. Planar heritage objects represent a significant category in computational reassembly, encompassing artifacts such as ancient maps, floor plans, frescoes, and inscribed tablets. A landmark project in this domain is Stanford’s Forma Urbis Romae project [262] (Figure 19 (Top-Left)), which aimed to reconstruct an ancient marble map of Rome. Despite scanning 1186 fragments (approximately 20% of the estimated total), the project faced significant challenges due to missing pieces. Their methodology combined multiple matching approaches: boundary incision matching using surface textural features, location prediction using map features, and multivariate clustering that considered characteristics like marble veining direction and fragment thickness [81].
@@
The Theran wall paintings project  [263, 264, 22, 265, 144, 266, 148] represents another significant body of work in planar heritage reconstruction, including a series of work at Princeton University on reconstruction [22], matching [264, 265, 144], and analysis of fragment boundary patterns for fragment pattern simulation [266]. Moreover, [148] presents a generic algorithm for reassembling wall paintings. One major difference between wall fragments and fragments of bones and thin-shells is that the number of wall fragments is much larger. In other words, it is important to develop computational solutions to fractured object reassembly. Another consequence of having a large number of fragments is that the global consistency constraint [144, 148] becomes important for multi-fragment matching.
For inscribed tablets, researchers have addressed the challenge of preserving both geometric and textual information. [267] developed methods for cuneiform tablet reconstruction considering both wedge impressions and surface geometry, while [268] introduced a system combining geometric and semantic analysis. More recent work [269] has used machine learning to analyze cuneiform characters for matching purposes.
Recent advances in this field include RePAIR dataset [270], which provides both 2D and 3D puzzles from collapsed frescoes, incorporating real-world challenges such as missing pieces and irregular fragment shapes. Related work on flat document restoration, such as papyrus reconstruction [271], has leveraged modern machine learning techniques, combining convolutional neural networks with traditional classifiers to address the unique challenges of document reassembly.
Stone fragments. The reassembly of stone fragments encompasses both individual artifacts like sculptures and large-scale architectural remains, presenting unique challenges due to material properties, scale, and frequent missing pieces. Early work [53] introduced geometric matching techniques specifically designed for fractured objects, combining patch-based features with global optimization to handle stone fragments effectively. For sculptural fragments, [97] developed methods based on original surface features, while [98] introduced robust scoring functions considering gap volume and contact surface properties. Recent work [95] proposed specialized methods for stone tool reassembly using contour points and flake surfaces.
In architectural reconstruction, the challenges scale significantly with fragment size and quantity. The Octagon monument project [51], employing techniques from [53] for pairwise matching and multi-piece assembly, demonstrated semi-automatic restoration using only 20% of original fragments. The Arch of Titus at the Circus Maximus in Rome reconstruction [210] leveraged architectural symmetry to handle extensive missing pieces. Similar approaches have been applied to other significant sites, including the Maritime Theatre [272] and Macedonian Tombs [273]. For specific architectural elements, such as column drums, [274] developed specialized pattern matching techniques. Recent advances [82] address eroded fragments by developing completion methods before matching. A key trend across both domains is the integration of computational methods with expert knowledge, particularly crucial when dealing with weathered surfaces or significant missing portions.
@@ 7.3 Fractured Object Reassembly in Science
There are many scientific applications for fractured object reassembly. One such example is docking of proteins and DNA Sequencing. In [275], which mitochondrial genes from the unicellular eukaryote Diplonema are systematically fragmented into small pieces that are encoded on separate chromosomes and individually transcribed. The reassembly problem is to concatenate them into contiguous messenger RNA molecules. Another biological application is protein docking [279, 276], where the goal is to find a rigid and non-rigid transformation of one 3D protein to interact with another protein.
Another application is in paleontology, in which we can reconstruct fossils from fragments precisely in 3D, to allow 3D analysis of the whole shape. A semi-automatic approach was introduced in [277] to reconstruct an extinct bird. It is desirable to develop automatic approaches for more accurate and faster reconstructions. [280] studied the virtual reconstruction of huminin crania to understand human evolution. [281] introduced an interactive approach in Blender to restore a dorsal rib and frontal bone of a small-bodied Jurassic-age armored dinosaur from Africa, The digital restoration engaged all modalities of deformation (translation, rotation, scaling, distortion) and reconstruction (fracture infilling, adding missing bone, surface smoothing).
@@
Table 3: Comparison of publicly available datasets for fractured object reassembly. Datasets are listed chronologically.
Datasets Objects/Groups Pieces per Object Total Pieces Breakdown Type Data Type Data Origin Ground Truth TU Wein [53] 7 6-30 101 Real 3D Artifacts Roughly Presious [284] 7 3-30 69 Real 3D Cathedrals Partial BreakingBad [149] 10474 2-100 8442044 Synthetic 3D - FantasticBreaks [285] 150 2 300 Real 3D Household FIRES [24] 15 7-18 123 Real 3D Pottery RePAIR [270] 18 2-44 1070 Real 2D/3D Frescoes Partial
@@ 7.4 Part Assembly
Part assembly refers to the process of assembling components that are decomposed in a semantically meaningful way, such as furniture components like table legs[201] and mechanical parts like screws[283]. Although closely related to fracture assembly problem, part assembly introduces several key differences. It generally assumes that each part has an independent semantic function and that each component is complete and functional on its own, unlike fractured pieces in fracture assembly. Moreover, part assembly allows for duplicated components, and solutions are often non-unique. Unlike fracture assembly, where local or global geometry provides significant clues, part assembly relies more on semantic or functional understanding of the individual parts and the assembled object. However, despite these differences, we can identify valuable connections between the two problems, and both fields can benefit from exploring the shared foundations of these challenges.
@@
Table 4: Benchmark performance of published methods on the Breaking Bad Dataset. All approaches were initially trained and tested on the everyday object subset. They are then fine-tuned and tested on the artifact subset (unless noted with b). Testing was conducted on objects containing 2-20 fragments. Methods are listed chronologically.
everyday artifact RMSE(R) MAE(R) RMSE(T) MAE(T) CD PA time RMSE(R) MAE(R) RMSE(T) MAE(T) CD PA degree degree ×$\times 10^{-2}$$\times 10^{-2}$$\times 10^{-3}$ ms degree degree ×$\times 10^{-2}$$\times 10^{-2}$$\times 10^{-3}$ Global [295] 80.7 69.7 14.8 11.8 14.6 21.8 23.7 83.8 71.8 16.6 13.8 19.0 13.3 LSTM [306] 84.2 72.7 16.2 12.7 15.8 19.4 29.9 84.6 73.1 16.8 14.0 21.5 11.7 DGL [212] 79.4 67.8 15.8 12.5 14.3 23.9 28.7 81.7 69.7 16.6 13.8 17.3 19.4 SE(3)-Equiv [146]c 79.3 - 16.9 - 28.5 8.4 129.9 - - - - - - Jigsaw [62] 42.3 36.3 10.7 8.7 13.3 57.3 1063.5 52.4 45.4 22.2 19.3 14.5 45.6 b PHFormer [154] 26.1 22.4 9.3 7.5 9.6 50.7 10.7 32.1 27.7 11.4 9.5 11.8 37.1 DiffAssemble [157] 73.3 - 14.8 - 27.5 - - - - - - - PuzzleFussion++ [145] 38.1 - 8.0 - 6.0 70.6 928.5 52.1 - 13.9 - 14.3 49.6 b PMTR [151] a 31.6 - 9.9 - 5.6 70.6 - 31.6 - 10.1 - 4.3 71.6 FragmentDiff [158] a 13.7 11.5 7.4 5.8 - 90.2 - 18.2 - 8.1 - - 82.3 b
  • a
    Reported model is trained and tested on the “volume-constrained” subset.
  • b
    Reported model is only trained on everyday subset and is not finetuned on the artifact subset.
  • c
    Original paper tested on 2-8 fragments; results for 2-20 fragments are cited from PuzzleFusion++.
Figure 22: Datasets used for deep learning research in fractured object reassembly. (Top-Left) Fantastic Breaks dataset [285] which combines both isolated fragments (top) and paired complete objects and their fragments (bottom). (Top-Right) A real dataset of fragments with ceramic material [24] is captured using the system shown here. (Bottom-Left) Synthetic dataset created by performing fracture simulation on a diverse set of objects [149]. (Bottom-Right) RePAIR dataset [270] which contains scans of a fresco at Pompeii archaeological park and ground truth provided by archaeologists.
8 Datasets and Open Source
We first discuss existing datasets in Section 8.1 and the evaluation metrics used in benchmarks in Section 8.2. We then discuss existing open-sourced methods in Section 8.3.
@@ 8.1 Datasets
Machine learning approaches have shown great advancements in the task of reassembly of fragments, but are highly dependent on large-scale datasets. However, the first data sets used in fragment reassembly research are primarily small-scale, often manually created [53, 307, 262, 284] or derived from automated semantic segmentation [302, 201].
@@
Several recent efforts have focused on scanning fragmented objects at scale. The first example is Fantastic breaks [285], which includes a small collection of objects with broken fragments and more objects that were broken into fragments by the authors (see Figure 22(Top-Left)). Both fragments and complete objects (if they exist) are scanned. [24] developed a system to batch scan fragments of objects with ceramic material (see Figure 22(Top-Right)). This dataset has ground-truth complete objects for evaluation. In addition, the scale of these two datasets makes it possible to develop deep neural networks for reassembly. However, the number of categories in these datasets is still small, presenting generalization issues in novel categories. The RePAIR dataset [270] (see Figure 22(Bottom-Right)) represents a new advance in real-world fragment assembly datasets, built from a collapsed fresco excavated from archaeological sites. It provides both 2D and 3D puzzles along with high-resolution texture information and expert archaeological annotations. It includes scenarios with missing pieces, highly irregular fragment shapes, and unknown matching relationships between pieces, which breaks a key assumption present in most existing object-level assembly datasets - that all pieces originate from a single, complete object.
@@
In addition to real data sets, Breaking Bad [149] is a synthetic data developed by running a simple fracture simulation procedure to generate fragments of shape collections that undergo large geometric and topological variations (see Figure 22(Bottom-Right)). This dataset is interesting for fractured object reassembly because the underlying objects belong to diverse object categories. For example, template-based reassembling techniques can be developed by learning a shape prior of the complete objects in this dataset. Just as the role of ShapeNet [308] in boosting 3D deep learning rich, we envision that research progress on algorithmic reassembly of fractured objects will be accelerated with Breaking Bad and similar datasets. However, the fragments are generated using non-physics based fracture simulation. The question is how the reassembly pipelines derived from such synthetic datasets generalize to real data. We present a comparison of publicly available datasets in Table 3.
Figure 23: Illustrations of future directions (Left) Representations and algorithms for matching complete template shapes with large geometric and topological variations and end-to-end learning of the entire reassembly pipeline. (Middle) Leveraging foundation models to solve sub-tasks such as segmenting fractured surfaces, matching fragments, predicting complete template shapes from input fragments, and matching complete template shapes and fragments. (Right) Learning to generate synthetic datasets of 3D fragments and doing so jointly with learning the reassembly network.
8.2 Evaluation Metrics
To quantitatively evaluate reassembly methods, the research community has adopted several standard metrics that measure both the accuracy of predicted transformations and the quality of the final assembly.
For rotation evaluation, two primary metrics are commonly used: the Mean Absolute Error (MAE) and Root Mean Square Error (RMSE). Given a predicted rotation $\tilde{R}$ and ground-truth rotation $R^{gt}$ represented in Euler angles, these metrics are defined as:
[…]
Similarly, for translation evaluation, given a predicted translation $\tilde{\boldsymbol{t}}$ and ground-truth translation $\boldsymbol{t}^{gt}$, the corresponding metrics are:
[…]
To evaluate the quality of the assembled result, two additional metrics are widely used. Chamfer Distance (CD) evaluates the overall assembly quality by measuring the bi-directional point cloud distance between the assembled result and the ground truth shape:
$\textup{CD}(X,Y)=\frac{1}{|X|}\sum_{x\in X}\min_{y\in Y}||x-y||_{2}+\frac{1}{|Y|}\sum_{y\in Y}\min_{x\in X}||y-x||_{2}$ (6)
Part Accuracy (PA) measures the percentage of correctly placed fragments. A fragment is considered correctly placed if its Chamfer distance to the ground truth is below a threshold (typically 0.01):
PA=number of fragments with CD<0.01total number of $\textup{PA}=\frac{\text{number of fragments with CD}<0.01}{\text{total number of fragments}}$ (7)
Finally, computational efficiency is measured through the total execution time of the algorithm, typically measured in milliseconds. This metric is particularly important for practical applications where real-time or near-real-time performance is desired.
For multi-piece reassembly tasks, the transformation errors (MAE and RMSE) and CD are typically computed as the mean error across all constituent pieces of each object. The overall performance of a method is then evaluated by averaging these object-level metrics across the entire test set. While these metrics provide a standardized way to compare different reassembly approaches, it is worth noting that they may not fully capture aspects such as physical validity or aesthetic quality of the final assembly.
8.3 Open Source